Twenty liters of a 20% acid solution is obtained by mixing 30% acid solution and a 5% acid solution. How many liters of each solution are needed to obtain the specified mixture?
Answer by jorel1380(3719)

20ltr.x.20=4 ltrs. pure acid
-----------------
m+n=20
.30m+.05n=4
30m+5n=400
5m+5n=100
25m=300
m=12
You need 12 ltrs. of 30% solution to 8 ltrs. of 5% solution..